OpenStudy (pinklion23):

Which ordered pairs in the form (x, y) are solutions to the equation 7x – 5y = 28? Choose all answers that are correct. A. (−6, −14) B. (−1, −7) C. (4, 10) D. (7, 9)

OpenStudy (pinklion23):

@dakid88

OpenStudy (dakid88):

well lets see.. 7x - 5y = 28 1) (-6,-14)...x = -6 and y = -14 now we sub 7(-6) - 5(-14) = 28 -42 + 70 = 28 28 = 28 (correct) --- this one is a solution 2) (-1, -7)...x = -1 and y = -7 now we sub 7(-1) - 5(-7) = 28 -7 + 35 = 28 28 = 28 (correct) -- this one works as well do you see how I am doing this ? Do you understand How to do it now ? I will leave the last two for you to figure out. Just reember, that at the end, if they do not equal, it is not a solution.

OpenStudy (pinklion23):

lol you copied and pasted that :P

OpenStudy (dakid88):

nope

OpenStudy (dakid88):

wrot it

OpenStudy (dakid88):

i mean typed it

OpenStudy (pinklion23):

sure well the answers are A and B found it :)

OpenStudy (dakid88):

oh well there u go

OpenStudy (dakid88):

hey how come u dont never text me anymore on gmail